# Hey Dashboard!: Supporting Voice, Text, and Pointing Modalities in Dashboard Onboarding

### v1

@@ 4 Implementation
For the natural language interaction, we use the OpenAI APIs: Chat Completions with gpt-4.1-mini for text [1] and the Realtime API [2] with gpt-4o-realtime-preview for voice I/O (with automatic speech transcription). All OpenAI calls are proxied through the Flask backend [30], which mints ephemeral Realtime session tokens so the browser never sees the primary API key.
@@
Diana cannot query the underlying dataset nor retrieve off-screen values. Therefore, it cannot provide data-derived answers or numeric values. It only uses the metadata and content available from the embedded dashboard (e.g., visual titles, layout bounds, and author-provided descriptions). Diana has no authoring control over the dashboard, and therefore, it cannot apply filters, or interact with visuals. When a user request requires information that is not available from the visible dashboard context, the system emits a structured “NOT IN DATA” outcome and redirects the user to the appropriate navigation or exploration steps.
5 Study Design and Methodology
@@
Our goal was to examine how users engage with Diana (DG2 and DG3: multimodal interaction and human-like onboarding) and how it affects the balance between guidance and user autonomy (DG5). We therefore adopted a qualitative, task-based study design combining think-aloud protocols with semi-structured interviews. We conducted the study in two phases: (1) a formative study (n=4), to refine the task design and the set of supported modalities, and (2) main interviews (n=6) with participants spanning novice to experienced dashboard users.
During the main sessions, participants completed a series of dashboard-oriented tasks of varying difficulty using Diana. Participants were free to choose the interaction modality (or modalities) per task and could ask follow-up questions on tasks in more detail when needed. Comparing real-world onboarding scenarios, we also provided a digital standard user guide for the dashboard as an optional reference but participants were not required to use it.
5.1 Participants
@@
We first did a formative study with four participants, out of which one was (A1) dashboard author who regularly onboards several colleagues and the other three (A2-A4) were domain experts and dashboard users with different level of visualization and dashboard expertise, as shown in Figure 3. We used the insights from the formative study to inform the task phrasing and overall improvement of the prototype for the main session. We do not include the formative study data in the main analysis.
@@
Both the formative study and main session participants were recruited from different university departments, such as Business Intelligence, Economics, Finance & Controlling, and Computer Science. For the main session, the average age of the participants was 36.7 years ($\sigma$ = 11.1) and their experience in their field of expertise ranged from 2 to 25 years. Out of the six main study participants, three were well-versed with dashboards (P2, P3, and P4), with P2 providing regular onboarding to many of his colleagues. The other three participants (P1, P5 and P6) were novices and had little to no experience with visualization dashboards. Each interview was conducted in person and individually.
@@
We also asked all these participants about their familiarity of AI tools and their experience with different modalities. These participants were chosen for two main reasons: (i) to understand how varied level of expertise (novice-expert) leads to differences in self-onboarding practices, and (ii) how does their experience with AI and multimodalities inform their preferences during the onboarding process.
@@
Figure 3: Participant demographics. Demographics of the formative study participants (A1-A4) and main study participants (P1-P6). Gender is encoded as male (M) and female (F). Visualization and dashboard experience is measured from novice to expert. AI interface experiences is noted for chat and voice interfaces.
5.2 Dataset and Task
@@
We used an example dataset from Microsoft Power BI [26] to create a custom visualization dashboard. The dashboard consisted of three key performance indicators (KPIs), one date filter, a funnel chart, a line chart, a clustered bar chart, and a map visualization. All visualizations, except the KPIs, were interactive. In addition to highlighting and filtering, a few visuals supported drill-down functionality, allowing users to explore information in greater depth.
@@
We leveraged this interactivity when designing the task set, which spanned three dimensions: task type, cardinality, and level of difficulty.
@@
  • Cardinality: This dimension captured the number of views a task required. As the dashboard consisted of multiple coordinated views, some tasks could only be solved by interacting with more than one visualization.
@@
For instance, an easy look-up task involving only a single view could be How do I figure out the scaling of the x-axis of the bar chart?. Meanwhile, a difficult task involving multiple views could be How can I figure out the revenue goal for Australia in the Services subcategory for the Proposal stage?. We designed ten such tasks based on our own onboarding experience and in discussion with the formative study participant A1. Figure 4 shows the task types and their difficulty levels. The task questions are provided in the supplementary material.
5.3 Study Procedure
Each session began with an introduction to the study goals and a request for informed consent regarding audio and screen recording. After receiving their consent, we asked participants to fill out a pre-questionnaire, containing a set of preliminary questions about their prior experience with visualizations, dashboards, and AI interfaces (summarized in Figure 3).
Following this, participants completed a short training session that lasted approximately five minutes on average. During the training, we introduced Diana and asked participants to try each interaction modality at least once, as well as to experiment with combining modalities. This was intended to ensure that participants felt confident in using Diana before starting the tasks with the dashboard. In addition to the modalities provided by the system, a dashboard user guide (as a digital document) was available as an optional resource throughout the study.
The main task phase of the study lasted around 25 minutes. Participants were free to choose whichever modalities they preferred for each task and could switch between them at any time. They could ask the system as many questions as they wanted until they were satisfied with an answer, or solve the tasks on their own without assistance. If a task could not be solved, participants were allowed to skip it and continue to the next one. After completing all tasks, participants were interviewed about their onboarding experience, comparisons to their usual onboarding practices, and their preferred modalities.
5.4 Data Collection and Analysis
To capture participants’ backgrounds and experiences, we collected data using the pre-questionnaire as well as follow-up interviews after the tasks focusing on onboarding practices, modality preferences, and overall reflections on the experience.
@@ 5.4 Data Collection and Analysis
For each task we recorded whether it was solved, the modality and any switches between modalities, the Diana features that participants engaged with, and their rationales for switching. In their post task interview, we collected qualitative data on their preferred modality, their perceived guidance and autonomy and their experience compared to their previous onboarding experiences. The time taken to complete a task was not taken into account as it depends heavily on how participants understood the task phrasing and their familiarity with various modalities.
6 Results
@@
In the following, we present the findings from the four formative study sessions and the six main study sessions.
Figure 4: Modalities and features used during onboarding. Filled boxes show the use of a modality or a feature. Empty boxes indicate that it was not used. Boxes are colored by participants ID: P1  , P2  , P3  , P4  , P5  , and P6  .
6.1 Formative Study Interviews
@@
We did a formative study with four participants (A1-A4) to refine both the design and implementation of Diana and the experimental procedure. The list of the participants included a Power BI specialist (A1), a regular Power BI user (A2), a dashboard author and occasional user (A3), and a daily Power BI user (A4). In the following, we describe their experience with dashboards, onboarding, and their impressions about the system.
@@
A1 regularly created onboarding materials for dashboards, typically using videos and bookmarks in Power BI. Their approach involved taking screenshots of dashboards, adding markers and pointers, and displaying them as a translucent overlay on the first page, which users can dismiss after viewing. They noted that this process is time-consuming and requires frequent updates as dashboards evolve. A1 also emphasized that dashboard features often contain subtle behaviors that are difficult to onboard effectively. For example, drill-down interactions usually confused users because the view changes immediately, and users may not recognize the update. During task performance, A1 suggested that visuals should be highlighted directly and that the AI should provide more specific guidance, while also conveying the purpose of the dashboard.
@@
A2 relied heavily on dashboards in Microsoft Fabric and typically used official documentation for onboarding. If documentation was insufficient, they turned to community forums, and only as a last resort, to AI systems. Having used voice interfaces for more than seven years, A2 expressed disappointment that these technologies have not met their early promises, preferring instead what they described as “ground truth” onboarding methods. They also noted that seeking help was particularly difficult for novices who may not yet know what to ask. A2 valued contextual menus that provided embedded information about a visualization. Based on this feedback, we included an additional option for onboarding: a user guide containing annotated images and instructions for how to read, interact with, and interpret the dashboard.
@@
A3 used Power BI dashboards infrequently. When joining her institution, she received a single 45-minute onboarding session covering relevant dashboards. For support, she typically experimented on her own or asked a mentor. During the tasks, A3 expected contextual menus to provide richer explanations. Although she did not normally use voice interfaces, in the study, she quickly shifted to voice to ask clarification questions—about abbreviations, interaction methods, and task instructions. She appreciated the multimodality of the onboarding, as she valued having options. While she preferred chatting via text over talking out loud, she emphasized the importance of having a choice. She also combined mouse-based selection with writing to find answers.
@@
A4 used dashboards daily and, like A3, had received only a single onboarding session. When encountering issues, she usually consults her mentor. Although she used Microsoft Co-Pilot for other tasks, she had not used it with dashboards. During the tasks, she favored chat for its step-by-step guidance, but also experimented with speech interaction. She appreciated that the Open AI’s Realtime API recognized her accent and changed to a different language accordingly, but suggested improvements such as synchronized animations, slower speech, and a transcription option. The user guide, by contrast, felt burdensome. A4 preferred using text and speech together: the chat allowed her to get written instructions while the voice output guided and confirmed her actions in real-time. She saw this combination as a way to save time for both herself and her mentor, reducing the need to ask experts for help.
@@
Our formative study highlighted four changes: (i) Visual highlighting is essential to direct attention and orient users within the dashboard, (ii) prompts must be instructional so that they can teach a user how to perfrom concrete actions (e.g., drill-down on a visualization), (iii) the information text box that appears with the contextual menu should have narrative style information with icons to enhance readability, and (iv) a ground-truth fallback is needed, such a user guide so that the users can use it, if they prefer documentations over Diana’s support.
6.2 Main Study
In the main study, we had six participants whose performance varied systematically with the task type, number of views required, and the difficulty level. They frequently switched modalities when tasks became complex, combining different modalities along with features such as lasso selections, radial menus, and visual highlights.
6.2.1 Easy Tasks
@@
Most participants solved easy lookup or hover tasks independently, relying little on Diana. Onboarding was mainly used for confirmation rather than discovery. For instance, P2 and P3 used speech interaction and employed the voice and visual highlights feedback to double-check their answers, while P5 used the chat interface to clarify how to solve one task. As P1 explained, she was “used to figuring things out on [her] own,” describing the multimodal help as “brand new… but pretty cool to have these options.” The learning effect was evident by the ninth task, which every participant solved without any onboarding, regardless of prior expertise.
6.2.2 Medium Tasks
@@
Medium-level tasks typically required interacting with two to three views and triggered frequent back-and-forth interactions with Diana and switching heavily between modalities. Participants P1 and P5 used the keyboard to chat with Diana. The text-based response, coupled with embedded visual highlighting of the relevant region, helped participants understand where to look at and execute the onboarding steps to solve the tasks (DG4). P1 remarked that “it is pretty clever that it doesn’t tell you the answer, because sometimes you write [or type] and you get the answer, and you don’t know how you actually got there! That’s pretty cool”.
@@
P2, P3, P4, and P6 used speech interaction, sometimes in combination with the mouse (DG2). The audio and visual feedback helped them orient themselves on the dashboard and find the right steps to get to their answers. P3 and P6 asked several follow-up questions to Diana, in order to interpret the question, engaging with at least two modalities at a time (DG2 and DG3). P2 also switched modalities, from keyboard to speech, to rephrase the question in a more natural-sounding way. P1 and P2 also used the mouse to interact with the radial menu to gain more information on the onboarding content. This illustrates how participants used Diana not just to retrieve answers, but also to scaffold their own orientation and reasoning in the dashboard.
6.2.3 Hard Tasks

### v2

@@ 4. Implementation
For the natural language interaction, we use the OpenAI APIs: Chat Completions with gpt-4.1-mini for text (OpenAI, 2023a) and the Realtime API (OpenAI, 2023b) with gpt-4o-realtime-preview for voice I/O (with automatic speech transcription). All OpenAI calls are proxied through the Flask backend (Pallets, 2023), which mints ephemeral Realtime session tokens so the browser never sees the primary API key.
@@
Diana cannot query the underlying dataset nor retrieve off-screen values. Therefore, it cannot provide data-derived answers or numeric values. It only uses the metadata and content available from the embedded dashboard (e.g., visual titles, layout bounds, and author-provided descriptions). Diana has no authoring control over the dashboard and therefore cannot apply filters or interact with visuals. When a user request requires information that is not available from the visible dashboard context, the system emits a structured “NOT IN DATA” outcome and redirects the user to the appropriate navigation or exploration steps.
5. Study Design and Methodology
Our goal was to examine how users engage with Diana (DG2 (multimodal) and DG3 (human-like)) and how it affects the balance between guidance and user autonomy (DG5 (guidance & autonomy)). To provide an appropriate baseline for comparison, we also included a condition where participants were not given access to Diana at all and could only use a standard static dashboard guide provided as a printed document.
@@ 5. Study Design and Methodology
We therefore adopted a mixed-methods approach, measuring task performance across both conditions and conducting qualitative analysis of think-aloud sessions and subsequent semi-structured interviews. We conducted the study in two phases: (1) a formative study (n=4), to refine the task design and the set of supported modalities, and (2) main interviews with 12 participants, including six participants using Diana (D1-D6) and six participants using a static dashboard guide (B1-B6). The participants ranged from novice to experienced dashboard users.
During the main sessions, participants completed a series of dashboard-oriented tasks of varying difficulty, either using Diana or the static dashboard guide.
All (n=12) participants received the static dashboard guide, but only half (D1-D6) had access to Diana in addition to the guide. Participants using Diana were free to choose any interaction modality (or modalities) per task and could ask follow-up questions on tasks to Diana if needed.
5.1. Participants
@@
We first conducted a formative study with four participants, out of which one was (A1) a dashboard author who regularly onboards several colleagues, and the other three (A2-A4) were domain experts and dashboard users with different levels of visualization and dashboard expertise, as shown in Figure 3. We used the insights from the formative study to inform the task phrasing and overall improvement of the prototype for the main session. We do not include the formative study data in the main analysis.
@@
Both the formative study and main session participants were recruited from different university departments, such as Business Intelligence, Economics, Finance & Controlling, and Computer Science. For the main session, the average age of the participants was 38.4 years ($\sigma$ = 10.2), and their experience in their field of expertise ranged from 2 to 25 years. The first six participants (D1-D6) received Diana, in addition to the standard user guide. Among them, three were well-versed with dashboards (D2, D3, and D4), with D2 providing regular onboarding to many of his colleagues. The other three participants (D1, D5 and D6) were novices and had little to no experience with visualization dashboards. Among the baseline participants (B1-B6) who received only the standard user guide, one had intermediate knowledge about the dashboards (B2), and one was an expert (B3). The other four participants (B1, B4, B5, and B6) were novices but had some experience with visualizations. Each interview was conducted in person and individually.
@@
We also asked all these participants about their familiarity with AI tools and their experience with different modalities. These participants were chosen for two main reasons: (i) to understand how varied level of expertise (novice to expert) leads to differences in self-onboarding practices, and (ii) how does experience with AI and multimodality inform user preferences during onboarding.
@@
Figure 3. Participant demographics. Overview of the formative study participants (A1-A4), Diana condition participants (D1-D6, and baseline condition participants (B1-B6). Gender is encoded as male (M) and female (F). Visualization and dashboard experience is measured from novice to expert. AI interface experiences are noted for chat and voice interfaces.
The demographics table lists each participant’s age, gender, visualization experience, dashboard experience, onboarding background, and prior AI interface experience. A pattern emerges in which younger participants, particularly those aged 25–30, tend to show higher visualization and dashboard experience as indicated by more filled bars, whereas older participants often exhibit less direct tool experience but may have engaged more with documentation or training materials. Chat-based AI experience is common across groups, but voice-based interaction appears less widespread. Differences also appear across participant groups: the A group frequently reports structured onboarding sources such as documentation or videos, the D group often has blended onboarding exposure, and the B group includes several participants with little or no formal onboarding experience.
5.2. Dataset and Task
@@
We used an example dataset from Microsoft Power BI (Microsoft Corporation, 2023b) to create a custom visualization dashboard. The dashboard (Figure 2) included three key performance indicators (KPIs), a date filter, a funnel chart, a line chart, a clustered bar chart, and a map visualization. All visualizations, except the KPIs, were interactive. In addition to highlighting and filtering, a few visuals supported drill-down, allowing users to explore information in greater depth.
@@
We leveraged this interactivity when designing the task set, which spanned task type, cardinality, and level of difficulty.
@@
  • Cardinality: This dimension captured the number of views a task required. As the dashboard consisted of multiple coordinated views, some tasks could only be solved by interacting with multiple visualizations.
@@
For instance, an easy look-up task involving only a single view (Figure 2) could be How do I figure out the scaling of the x-axis of the bar chart? Meanwhile, a difficult task involving multiple views could be How can I figure out the revenue goal for Australia in the Services subcategory for the Proposal stage? We designed ten such tasks based on our own onboarding experience and in discussion with the formative study participant A1. Figure 4 shows the tasks, their types, and their difficulty levels.
5.3. Study Procedure
Each session began with an introduction to the study goals and a request for informed consent regarding audio and screen recording. After receiving their consent, we asked participants to complete a pre-questionnaire containing preliminary questions about their prior experience with visualizations, dashboards, and AI interfaces (summarized in Figure 3).
Diana
After the pre-questionnaire, the participants (D1-D6) completed a short training session lasting approximately five minutes. During the training, we introduced Diana on the dashboard and asked participants to try each interaction modality at least once and to experiment with combining modalities. This was intended to ensure that participants felt confident in using Diana before starting the tasks with the dashboard. In addition to the modalities provided by the system, a dashboard user guide (as a digital document) was available throughout the study.
The task phase of this study lasted around 25 minutes. Participants were free to choose whichever modalities they preferred for each task and could switch between them at any time. They could ask the system as many questions as they wanted until they were satisfied, or solve the tasks on their own without assistance.
Static Dashboard Guide
The static dashboard guide included text from the Microsoft Power BI sample dashboard webpage (Microsoft Corporation, 2023b), which provides a brief overview of the visuals and how to interact with them. In addition, drawing on our prior onboarding experience, we included a section that explains each visual, including its name, title, and an accompanying figure. We also added the dashboard’s purpose, a brief overview of available interactions, and general usage tips, resulting in a guide totaling 3.5 pages. B1 to B6 received the standard user guide and were given time to review it and use it throughout the study. They could choose whether to enter the task phase immediately or go through the documentation first. The task phase of this study lasted around 30 minutes.
In both study scenarios, if a task could not be solved, participants were allowed to skip it and continue to the next one. After completing all tasks, participants were interviewed about their onboarding experience, comparisons with their usual onboarding practices, and preferred modalities (only for participants in the Diana condition).
5.4. Data Collection and Analysis
To capture participants’ backgrounds and experiences, we collected data using the pre-questionnaire and follow-up interviews after the tasks, focusing on onboarding practices, modality preferences, and overall reflections on the experience.
@@ 5.4. Data Collection and Analysis
For each task we recorded whether participants successfully completed it and whether they used the provided onboarding material or assistance. For participants using Diana, we also noted the modality, any switches between modalities, the features participants engaged with, and their rationales for switching.
@@
After each task, we conducted brief interviews with the participants to gather qualitative feedback on their perceived level of guidance and autonomy. For the Diana group, we also asked about their preferred modalities and their experiences compared to prior onboarding approaches. The time required to complete a task was not taken into account, as it depends heavily on participants’ understanding of the tasks and their familiarity with the various modalities.
6. Results
@@
In the following, we present the findings from the four formative study sessions and Diana and static dashboard guide sessions.
6.1. Formative Study Interviews
@@
We conducted a formative study with four participants (A1-A4) to refine both the design and implementation of Diana and the experimental procedure. The list of participants included a Power BI specialist (A1), a regular Power BI user (A2), a dashboard author and occasional user (A3), and a daily Power BI user (A4). In the following, we describe their experience with dashboards and onboarding, and their impressions of the system.
@@
A1 regularly created onboarding materials for dashboards, typically using videos and bookmarks in Power BI. Their approach involved taking screenshots of dashboards, adding markers and pointers, and displaying them as a translucent overlay on the first page, which users can dismiss after viewing. They noted that this process is time-consuming and requires frequent updates as dashboards evolve. A1 also emphasized that dashboard features often contain subtle behaviors that are difficult to onboard effectively. For example, drill-down interactions usually confused users because the view changes immediately, and users may not recognize the update. During task performance, A1 suggested that visuals should be highlighted directly and that the AI should provide more specific guidance, while also conveying the dashboard’s purpose.
@@
A2 relied heavily on dashboards in Microsoft Fabric and typically used official documentation for onboarding. If documentation was insufficient, they turned to community forums, and only as a last resort, to AI systems. Having used voice interfaces for more than seven years, A2 expressed disappointment that these technologies have not met their early promises, preferring instead what they described as “ground truth” onboarding methods. They also noted that seeking help was particularly difficult for novices who may not yet know what to ask. A2 valued contextual menus that provided embedded information about a chart. Based on this feedback, we added an additional onboarding option: a user guide with annotated images and instructions for reading, interacting with, and interpreting the dashboard.
@@
A3 used Power BI dashboards infrequently. When joining her institution, she received a single 45-minute onboarding session covering relevant dashboards. For support, she typically experimented on her own or asked a mentor. During the tasks, A3 expected contextual menus to provide richer explanations. Although she did not normally use voice interfaces, during the study, she quickly shifted to voice to ask clarification questions—about abbreviations, interaction methods, and task instructions. She appreciated the multimodality of the onboarding, as she valued having options. While she preferred chatting via text over talking out loud, she emphasized the importance of having a choice. She also combined mouse-based selection with writing to find answers.
@@
A4 used dashboards daily and, like A3, had received only a single onboarding session. When encountering issues, she usually consults her mentor. Although she used Microsoft Co-Pilot for other tasks, she had not used it with dashboards. During the tasks, she favored chat for its step-by-step guidance, but also experimented with speech interaction. She appreciated that the Open AI’s Realtime API recognized her accent and changed to a different language accordingly, but suggested improvements such as synchronized animations, slower speech, and a transcription option. The user guide, by contrast, felt burdensome. A4 preferred using text and speech together: the chat allowed her to get written instructions while the voice output guided and confirmed her actions in real time. She saw this combination as a way to save time for both herself and her mentor, reducing the need to ask experts for help.
@@
Our formative study highlighted four changes: (i) Visual highlighting is essential to direct attention and orient users within the dashboard, (ii) prompts must be instructional so that they can teach a user how to perform concrete actions (e.g., drill-down on a visualization), (iii) the information text box that appears with the contextual menu should have narrative style information with icons to enhance readability, and (iv) a ground-truth fallback is needed, such as a user guide so that the users can use it, if they prefer documentations over Diana’s support.
6.2. Diana Condition
In both studies, participants completed the same set of dashboard-related tasks under two conditions: six participants used Diana in addition to the standard user guide, and six used the standard user guide only. Their performance varied by task type, number of views involved, and task difficulty (shown in Figure 4). The Diana group frequently switched modalities as tasks grew more complex, whereas baseline participants relied solely on textual instructions and standard dashboard interactions. We describe the study results below, first for Diana and then for the static dashboard guide.
Figure 4. Order of modalities and features used by Diana participants during tasks. Tasks are ordered by the order in which they were given to the participants. They are colored by their difficulty level: easy, medium, and difficult. Modalities are represented through colors and icons: mouse , keyboard , and speech-based interactions. Visual highlights and radial menu are also specified. The dashboard guide was not used by anyone and is not represented in the figure. An empty cell indicates that the participant solved the task on their own without the use of Diana.
This table presents ten tasks along with their difficulty, task type, and the interaction modalities used by each Diana participant. The overall pattern shows that Hard tasks consistently require more complex interaction strategies, with participants combining modalities such as hovering, clicking, typing, or voice commands to complete them. Easy tasks tend to rely on a single, simple interaction mode, usually involving only mouse actions. Medium tasks fall in between, with variability in whether participants resort to voice input or keyboard actions. Some participants repeatedly employ a wide range of modalities, while others rely on a narrower set, resulting in a visible correlation between task complexity and the breadth of interaction required.
6.2.1. Easy Tasks
@@
Most participants solved easy lookup or hover tasks independently, relying little on Diana. Onboarding was mainly used for confirmation rather than discovery. For instance, Task 4, which required looking up a catgeorycategory value in the bar chart, D1 solved it on her own as she is “used to figuring things out on [her] own,” describing the multimodal help as “brand new… but pretty cool to have these options.” (shown in Figure 4). Meanwhile, D2 and D3 used speech interaction and employed the voice and visual highlights feedback to double-check their answers, while D5 used the chat interface to clarify how to solve that task.
The learning effect was evident by Task 9 (shown as an empty row in Figure 4), which required selecting a region on the map and reading the relevant value on hover or from one of the Key Performance Indicators (KPIs). Every participant solved this task without any onboarding, regardless of prior expertise.
6.2.2. Medium Tasks
@@
Medium-level tasks typically required interacting with two to three views and triggered frequent back-and-forth interactions with Diana and switching heavily between modalities. Participants D1 and D5 used the keyboard to chat with Diana for most of the tasks. The text-based response, coupled with embedded visual highlighting of the relevant region, helped participants understand where to look and how to execute the onboarding steps to complete the tasks (DG4 (orientation)). While solving Task 1, which required a lookup and interpretation within the bar chart, D1 remarked that “it is pretty clever that it doesn’t tell you the answer, because sometimes you write [or type] and you get the answer, and you don’t know how you actually got there! That’s pretty cool.”
@@
For almost all medium tasks, D2, D3, D4, and D6 used speech interaction, sometimes in combination with the mouse (DG2 (multimodal)). For instance, in Task 2, which required locating and interpreting the answer in the line chart, D2 first used the chat interface to ask the question. They then used the visual highlights feature within the chat interface to orient themselves in the dashboard and then switched to the voice interface to ask the question again. Since Diana does not provide direct answers, they then used the radial menu to discover the insight needed to answer it. The audio and visual feedback also helped D3, D4, and D6 orient themselves on the dashboard and find the right steps to get to their answers. D2 also switched modalities, from keyboard to speech, to rephrase the question in a more natural-sounding way. D1 and D2 also used the mouse to navigate the radial menu to learn more about the onboarding content. This illustrates how participants used Diana not only to retrieve answers, but also to scaffold their own orientation and reasoning in the dashboard.
6.2.3. Hard Tasks